A theater group made appearances in two cities. The hotel charge before tax in the second city was
$1500
lower than in the first. The tax in the first city was
8.5%
, and the tax in the second city was
8%
. The total hotel tax paid for the two cities was
$457.50
. How much was the hotel charge in each city before tax?

Answer by ankor@dixie-net.com(22740)

A theater group made appearances in two cities.
The hotel charge before tax in the second city was $1500 lower than in the first.
The tax in the first city was 8.5%, and the tax in the second city was 8%.
The total hotel tax paid for the two cities was $457.50.
How much was the hotel charge in each city before tax?
:
let a = hotel charge for the 1st city
The 2nd city was $1500 lower, therefore
(a-1500) = hotel charge for 2nd city
:
1st Hotel tax + 2nd Hotel tax = 457.50
.085a + .08(a-1500) = 457.50
.085a + .08a - 120 = 457.50
.165a = 457.50 + 120
.165a = 577.50
a = 577.50/.165
a = $3500 charge for the 1st hotel before tax
and
3500 - 1500 = $2000 charge for the 2nd hotel
:
:
Confirm this solution, find the actual tax on each, find their sum.
.085(3500) = 297.50
.08(2000) = 160.00
---------------------
Total tax: 457.50